OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the first equation. x + 6y + z = 2 x – 3y + 2z = –1 x + 5y + 3z = –4 11y + 4z = 6 2y + 5z = 3 y – 2z = 6 –8y – z = 3 3y + 3z = –3 11y + 4z = –6 –9y + z = –3 –y + 2z = –6

OpenStudy (amorfide):

x=-6y-z+2 sub into x in each equation x – 3y + 2z = –1 x + 5y + 3z = –4 these will go to -6y-z+2– 3y + 2z = –1 -6y-z+2+ 5y + 3z = –4 these simplify too –9y + z = –3 –y + 2z = –6

OpenStudy (anonymous):

thanks